OpenStudy (anonymous):

Function value is sinα = 0 and the constraint is secα =-1. I don't know how to find all six trig functions

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(\theta)=0}} \\ \quad \\ \quad \\ sec(\theta)=-1\implies \cfrac{1}{cos(\theta)}=-1\implies \cfrac{1}{-1}=cos(\theta)\implies {\color{blue}{ -1=cos(\theta)}}\) any ideas where that might be?